Let f(x, y) ={ y³ /x² + y² if (x, y) ≠(0, 0) and 0 otherwise


Show that f is continuous but not differentiable at (0, 0).


1
Expert's answer
2022-01-25T12:56:53-0500

To show that the function


"f(x,y)=\begin{cases} \frac{y^3}{x^2+y^2}~~~~(x,y) \neq (0,0)\\
0~~~~~~~~~~(x,y)=(0,0)\end{cases}"

is continuous at "(0,0)" , we must show that "\lim_{9x,y)\to (0,0) }f(x,y)=f(0,0)"

By definition, "f(0,0)=0".

Let "y=mx" . Thus,

"\lim_{(x,y)\to (0,0)} f(x,y)=\lim_{x \to 0}f(x,mx)=\lim_{x\to 0} \frac{(mx)^3}{x^2+(mx)^2}=\lim_{x \to 0} \frac{m^3x}{1+m^2}=0"

Thus the limit is 0 which is equal to "f(0,0)"

Hence, "f(x,y)" is continuous at "(0,0)"


Now, to show that it is not differentiable at (0,0)

Differentiating with respect to x yields:


"f_x(x,y)=\frac{-2xy^3}{(x^2+y^2)^2}"

Now, "f_x(0,0)=0" , but

Consider the limits along the straight line "y=mx"

"\lim_{(x,y)\to (0,0)}f_x(x,y)=\lim_{x\to 0} \frac{-2x(mx)^3}{(x^2+(mx)^2)^2}= \lim_{x\to 0}\frac{-2m}{(1+m^2)^2}=\frac{-2m}{(1+m^2)^2}"


Since the limit along a straight line depends on the slope of the line. Therefore, the two-variable limit does not exist at (0, 0). Since the limit does not exist, this shows that the partial derivative of the function with respect to x exist but is not continuous.This shows that the function is not differentiable.